# Research Directions and Modeling Guidelines for Industrial Internet of Things Applications

### v1

@@ Abstract
The Industrial Internet of Things (IIoT) paradigm has emerged as a transformative force, revolutionizing industrial processes by integrating advanced wireless technologies into traditional procedures to enhance their efficiency. The importance of this paradigm shift has produced a massive, yet heterogeneous, proliferation of scientific contributions. However, these works lack a standardized and cohesive characterization of the IIoT framework coming from different entities, like the 3rd Generation Partnership Project (3GPP) or the 5G Alliance for Connected Industries and Automation (5G-ACIA), resulting in divergent perspectives and potentially hindering interoperability. To bridge this gap, this article offers a unified characterization of (i) the main IIoT application domains, (ii) their respective requirements, (iii) the principal technological gaps existing in the current literature, and, most importantly, (iv) we propose a systematic approach for assessing and addressing the identified research challenges. Therefore, this article serves as a roadmap for future research endeavors, promoting a unified vision of the IIoT paradigm and fostering collaborative efforts to advance the field.
@@ I Introduction
The advent of the fourth industrial revolution, commonly referred to as Industry 4.0, is revolutionizing the manufacturing landscape through the integration of Information and Communications Technologies (ICT) [1]. This paradigm shift is characterized by the digitalization of production processes, and the adoption of cyber-physical systems that seamlessly blend physical assets with digital technologies. In this regard, the Industrial Internet of Things (IIoT) emerges as a pivotal facilitator within the Industry 4.0 framework, driving the Industry 4.0 paradigm from traditional, rigid industrial processes to interconnected, adaptive, and more efficient systems, by envisioning the adoption of wireless communications in traditional industrial processes to augment their efficiency and security [2]. Specifically, industrial assets, such as robotic arms, pumps, valves, pistons, etc., are equipped with wireless devices embedding sensors and actuators to facilitate a myriad of novel applications, including analytics, diagnostics, monitoring, asset tracking, as well as various control use-cases encompassing process, regulatory, supervisory, and safety control.
In light of this, factories pivot away from the traditional reliance on extensive wired communication technologies, embracing pioneering wireless solutions that unlock unprecedented levels of flexibility and adaptability. However, the performance offered by current wireless technologies proves adequate for a limited range of applications, such as simple non-real-time monitoring. Yet, it falls short in supporting more demanding processes, like digital twins, motion controls, etc. Moreover, due to the diverse requirements inherent in various industrial operations coexisting within the same facility, contemporary wireless technologies are unable to comprehensively address all needs.
Consequently, the literature has proposed several technological advancements to fill this gap. From the standardization viewpoint, the 5th Generation (5G) New Radio (NR) standard was thought to overcome these limitations and unleash the real potential of the Industry 4.0 framework. Indeed, 5G networks are flexible systems capable of adapting to different verticals and serving a plethora of services. In particular, the 3rd Generation Partnership Project (3GPP), through its recent releases (i.e., Release 16 and beyond), has brought specific features to support Ultra-Reliable Low-Latency Communications (URLLC), along with the Non-Public Network (NPN) paradigm, to further enhance network performance and security without the support of service providers. Although this evolutional path has given a massive boost to the effective realization of the fourth industrial revolution, some industrial use cases still pose stringent requirements that cannot be fulfilled even by private 5G networks, paving the way for new 6th Generation (6G)-oriented research trends. In this regard, the academic world is already intensively shaping the evolution towards 6G networks, by proposing important paradigm shifts, like the utilization of the THz band, the introduction of radar capabilities in next-generation Base Stations (BSs), the interplay between Artificial Intelligence (AI) and radio communication technologies, or the advent of quantum technologies.
@@
Despite this ambitious evolutionary trend, the majority of contributions to the literature on 6G networks do not account for the specificities of industrial environments, potentially limiting their practical value in that context. For example, channel models should account for the unique dynamicity and density of obstacles prevalent in industrial environments, where, e.g., Non-Line-of-Sight (NLoS) links are widespread, and, at the same time, network protocols should account for a heterogeneous set of challenging offered traffics that also coexist on the same factory floor.
This may be attributed to the lack of a standardized and cohesive characterization of the 6G-oriented IIoT framework, since different entities, such as 3GPP and 5G Alliance for Connected Industries and Automation (5G-ACIA), propose their own perspective and nomenclature for IIoT application domains and Key Performance Indicators (KPIs).
@@
    Provide a unified vision of the main characteristics of industrial contexts, in terms of nomenclature, Application Domains (ADs), KPIs, and traffic types;
@@
    Characterize the IIoT requirements for 6G-oriented communications by performing a meticulous examination of the literature, encompassing white papers and standards, and by unifying their divergent aspects (e.g., consolidating different terminologies that refer to the same KPI);
@@
    Identify the principal technological gaps existing in the current literature and present our perspective on the emerging research themes across various layers of the ISO-OSI standard;
@@
    Propose guidelines for assessing and addressing the identified research challenges to promote reproducibility and compatibility among different 6G-oriented scientific contributions.
@@ II IIoT Application Domains
In this section, we establish a unified nomenclature and description of application domains for the IIoT. It is worth highlighting that we adopt naming conventions from existing literature, without suggesting any modification or addition. Our goal is to clarify these terms and provide a comprehensive collection of their meanings to enhance understanding and facilitate consistent usage.
@@ III IIoT Characterization
Communication service Survival Time ST Maximum amount of time that the communication service could fail to meet the application message delay requirement before an application layer failure occurs such that the communication service is considered to be in an unavailable status ms
@@
Communication Service Reliability /// Reliabity CSR Mean value of how long the communication service is available before it becomes unavailable year
@@
We then hereby describe the main KPIs, the typical traffic types, and requirements to be adopted in the aforementioned 6G-oriented IIoT paradigm.
@@ III-B Traffic types
  • Non-Deterministic Communication (NDC) encompasses all traffic types other than periodic/aperiodic deterministic communication. In this case, there are no requirements either in terms of E2EL or CSA. Similar to the ADC, the traffic is defined through SBR.
@@ IV Research Directions and Modelling Guidelines
In this section, we identify the key technological gaps in the current literature and offer our perspective on how emerging research fields could address these gaps to fulfill the 6G-oriented IIoT requirements.
@@ IV-A Research Directions
Open challenges: The high density of obstacles (e.g., machinery) in indoor industrial environments can hinder signal propagation causing reflections, scattering, and attenuation. The development of ad-hoc network layout design paradigms and innovative signal processing techniques is crucial for guaranteeing high communication and sensing performance.
Recently, Reconfigurable Intelligent Surfaces (RIS) have emerged as a means to optimize communication links and improve sensing coverage, nevertheless introducing an additional layer of complexity to network management [8]. Moreover, localizing and tracking objects transitioning from outdoor to indoor scenarios (and vice versa), as well as combining multi-technologies indoor localization techniques to meet the desired requirements, are still open problems.
@@
Vehicular Communications. Mobile robots, particularly relevant to AD 4, will play a significant role in modern industries by supporting production processes. In particular, UAVs can act as mobile base stations or relays, establishing network connectivity even in areas with challenging terrain or temporary infrastructure needs, or during emergencies. To enable mobile robot operations, a robust backbone network ensuring efficient and dependable communication within the robots and the control systems is required. In this context, Vehicle-to-Everything (V2X) emerging standards facilitate communication among mobile robots, infrastructure, and other devices, improving safety and traffic management. In particular, the possibility to have sidelinks that allow devices within proximity to communicate efficiently, forming ad-hoc networks in environments with limited infrastructure or in scenarios requiring rapid data exchange, is of paramount importance.
@@
Cryptography and Security. The last point we want to highlight is the need for robust encryption algorithms and authentication mechanisms in all ADs to ensure data integrity and confidentiality within industrial wireless networks, safeguarding sensitive information [11]. The necessity to adhere to stringent E2EL requirements limits the use of current complex cryptographic methods and advanced cross-layer authentication architectures, creating the need for the development of innovative encryption and authentication techniques.
@@ IV-B Modeling Guidelines
In the following, we present modeling guidelines for industrial environments that adhere to the unified IIoT framework proposed by this work. This approach holds significant importance in establishing a unified framework for scientific endeavors aimed at addressing the research gaps previously highlighted. Indeed, the applicability of existing literature contributions to IIoT environments strongly depends on the level of realism of their underlying assumptions and characterizations. Moreover, the comparison between different solutions is possible only when they rely on common and unified models.
@@
Industrial environments possess distinctive characteristics that render wireless propagation inherently unique and distinguishable from other radio channels. Typically, industrial premises exhibit larger dimensions and/or greater height compared to office or residential buildings. Furthermore, the size, density, and spatial distribution of industrial equipment can vary significantly from one case to another but, on average, constitute a highly dense scatterer environment. In light of these peculiarities, it becomes imperative to tailor channel models to suit industrial scenarios adequately. For instance, the 3GPP has proposed a path loss model specifically designed for indoor factories [^2] It is noteworthy that the term “factory” in 3GPP corresponds to the term “building” used in this article[14]. However, more detailed wideband approaches can be obtained by leveraging (i) ray-tracing tools, as propagation in factories is likely to be more site-specific than propagation in usual residential or office environments, (ii) channel measurements, as a proper calibration method for ray-tracing tools, or (iii) the utilization of AI. Indeed, the latter is recently gaining attention to discern the intricate relationship between common propagation markers and the key properties of industrial environments [15].
@@ V Conclusions
In this article, we presented a unified vision of the IIoT framework, a key enabler of the emerging fourth industrial revolution. Our focus included harmonizing diverse contributions from entities like 3GPP and 5G-ACIA in terms of nomenclature, ADs, KPIs, and traffic types. After a careful literature analysis, we also characterize the main IIoT requirements by harmonizing heterogeneous terminologies. Most importantly, we offered an overview of the main research topics and challenges within the IIoT paradigm according to the 6G roadmap, highlighting the potential for innovative industrial applications. Finally, we proposed modeling guidelines as a foundational framework for future research endeavors, addressing the identified research directions towards next-generation IIoT.

### v2

@@ Abstract
The Industrial Internet of Things (IIoT) paradigm has emerged as a transformative force, revolutionizing industrial processes by integrating advanced wireless technologies into traditional procedures to enhance their efficiency. The importance of this paradigm shift has produced a massive, yet heterogeneous, proliferation of scientific contributions. However, these works lack a standardized and cohesive characterization of the IIoT framework coming from different entities, like the 3rd Generation Partnership Project (3GPP) or the 5G Alliance for Connected Industries and Automation (5G-ACIA), resulting in divergent perspectives and potentially hindering interoperability. To bridge this gap, this article offers a characterization of (i) the main IIoT application domains, (ii) their respective requirements, (iii) the principal technological gaps existing in the current literature, and, most importantly, (iv) we propose a systematic approach for assessing and addressing the identified research challenges. Therefore, this article serves as a roadmap for future research endeavors, promoting a unified vision of the IIoT paradigm and fostering collaborative efforts to advance the field.
@@ I Introduction
The advent of the fourth industrial revolution, commonly referred to as Industry 4.0, is revolutionizing the manufacturing landscape through the integration of Information and Communications Technologies (ICT) [1]. This paradigm shift is characterized by the digitalization of production processes, and the adoption of cyber-physical systems that seamlessly blend physical assets with digital technologies. In this regard, the Industrial Internet of Things (IIoT) emerges as a pivotal facilitator within the Industry 4.0 framework, driving the Industry 4.0 paradigm from traditional, rigid industrial processes to interconnected, adaptive, and more efficient systems, by envisioning the adoption of wireless communications in traditional industrial processes to augment their efficiency and security [2]. Industrial assets, such as robotic arms or valves, are equipped with wireless devices to support applications like analytics, monitoring, and control (detailed descriptions are provided in [3]).
In light of this, factories pivot away from the traditional reliance on extensive wired communication technologies, embracing wireless solutions that unlock unprecedented levels of flexibility and adaptability [1]. Wi-Fi©, for instance, delivers high performance and user-friendly operation at a low cost but lacks energy efficiency, making it impractical for large-scale deployments involving battery-powered devices. LoRa©, on the other hand, offers long-range, low-power connectivity, making it well-suited for non-real-time monitoring applications, though its network throughput is severely limited. Similarly, ZigBee© and Bluetooth© are effective for short-range, low-power communication, ideal for specific applications like asset tracking or environmental monitoring, but they also struggle with limited data throughput [1]. 5th Generation (5G) New Radio (NR) also emerged as a viable candidate for industrial communications thanks to key features such as Ultra-Reliable Low-Latency Communications (URLLC), the Non-Public Network (NPN) paradigm, and integration with the widely adopted Time-Sensitive Networking (TSN) standard [4, 2]; as of today, however, it remains both complex and expensive to deploy and maintain. Overall, the real-world adoption of these technologies remains limited because they are unable to comprehensively address all the diverse, stringent, and co-located requirements inherent in industrial environments.
This limitation highlights the necessity for further innovation and sets the stage for the development of next-generation mobile radio networks. In this regard, the academic world is already intensively shaping such evolution, by proposing important paradigm shifts, like the use of the THz band, radar in next-generation Base Stations (BSs), or Artificial Intelligence (AI)-radio communication interplay.
@@
Despite this ambitious evolutionary trend, the majority of contributions to the literature do not account for the specificities of industrial environments (e.g., heterogeneous traffic), potentially limiting their practical value in that context.
This may be attributed to the lack of a standardized and cohesive characterization of the next-generation IIoT framework, since different entities, such as 3rd Generation Partnership Project (3GPP) and 5G Alliance for Connected Industries and Automation (5G-ACIA), propose their own perspective and nomenclature for IIoT application domains and Key Performance Indicators (KPIs).
@@
    Provide a holistic description of the main characteristics of industrial contexts, in terms of nomenclature, Application Domains (ADs), KPIs, and traffic types;
@@
    Characterize the IIoT requirements for future wireless communications by performing a meticulous examination of the literature, encompassing white papers and standards, and by unifying their divergent aspects (e.g., consolidating different terminologies that refer to the same KPI);
@@
    Identify the principal technological gaps in the current literature and present our perspective on the emerging research themes across various layers of the ISO-OSI standard;
@@
    Propose guidelines for assessing and addressing the identified research challenges to promote reproducibility and compatibility among different scientific contributions.
@@ II IIoT Application Domains
In this section, we clarify the different and sometimes contrasting nomenclatures used in the literature to describe application domains for the IIoT. Our goal, rather than proposing modifications or additions to existing naming conventions, is to provide a clearer understanding of these terms and to facilitate their consistent usage across various contexts.
@@ III IIoT Characterization
Communication service Survival Time ST Maximum time interval over which the communication service can fail to meet the application message delay requirement before an application layer failure occurs, leading to the communication service being considered as unavailable ms
@@
Communication Service Reliability /// Reliabity CSR Mean value of time intervals between consecutive unavailability events for the communication service year
@@
We then hereby describe the main KPIs, the typical traffic types, and requirements to be adopted in the IIoT paradigm.
@@ III-B Traffic types
  • Non-Deterministic Communication (NDC) encompasses all traffic types other than periodic/aperiodic deterministic communication. In this case, there are no requirements either in terms of E2EL or CSA. As in the case of ADC, the traffic is defined through SBR.
@@ IV Research Directions and Modelling Guidelines
In this section, we identify the key technological gaps in the current literature and offer our perspective on how emerging research fields could address these gaps to fulfill the next-generation IIoT requirements.
@@ IV-A Research Directions
Open challenges: The high density of obstacles (e.g., machinery) in indoor industrial environments can hinder signal propagation causing reflections, scattering, and attenuation. The development of ad-hoc network layout design paradigms and innovative signal processing techniques is crucial for guaranteeing high communication and sensing performance. Recently, Reconfigurable Intelligent Surfaces (RIS) have emerged as a means to optimize communication links and improve sensing coverage, nevertheless introducing an additional layer of complexity to network management [8]. Moreover, localizing and tracking objects transitioning from outdoor to indoor scenarios (and vice versa), as well as combining multi-technologies indoor localization techniques to meet the desired requirements, are still open problems.
@@
Vehicular Communications. Mobile robots, particularly relevant to AD 4, will play a significant role in modern industries by supporting production processes. In particular, UAVs can act as mobile base stations or relays, establishing network connectivity even in areas with challenging terrain or temporary infrastructure needs, or during emergencies. To enable mobile robot operations, a robust backbone network ensuring efficient and dependable communication within the robots and the control systems is required. In this context, Vehicle-to-Everything (V2X) emerging standards facilitate communication among mobile robots, infrastructure, and other devices, improving safety and traffic management. In particular, the possibility to have sidelinks that allow devices in spatial proximity to communicate efficiently, forming ad-hoc networks in environments with limited infrastructure or in scenarios requiring rapid data exchange, is of paramount importance.
@@
Security and Safety. The last point we want to highlight is the need for robust encryption algorithms and authentication mechanisms in all ADs to ensure data integrity and confidentiality within industrial wireless networks, safeguarding sensitive information [11]. Equally critical, especially in industrial environments involving human-machine interaction, is functional safety, which focuses on preventing hazardous failures by ensuring reliable operation of safety-critical systems, even under fault conditions (see Fig. 2). From a communication standpoint, achieving functional safety requires dedicated safety communication protocols capable of verifying the integrity, authenticity, and timeliness of transmitted data. Addressing these challenges necessitates integrating safety communication protocols with encryption and authentication techniques to meet stringent E2EL requirements.
@@
Figure 2: Experimental PoC for AD 4. The focus of this PoC is on mitigating risks associated with an Autonomous Mobile Robot (AMR), which autonomously navigates industrial spaces and transports hazardous liquids. Although the AMR is equipped with emergency systems (e.g., proximity sensors and lasers) that trigger abrupt stops to avoid collisions, sudden braking can lead to liquid spills, posing a threat to human operators. The block diagram of the considered solution is illustrated in a). The AMR is equipped with a motion sensor, and the collected data are sent to a local 5G board powered by a power bank. The board transmits the data to a local server via a private 5G NPN, consisting of a dedicated Radio Access Network (RAN) and Core Network (CN). The local server, housed within the same rack as the 5G CN, hosts AI algorithms that process real-time motion data to predict potential spills. An example of a prediction is shown in b), where the x-axis acceleration (Ax in the legend) is plotted over time, with time discretized based on the 10 ms acquisition periodicity of the motion sensor. The yellow vertical bar indicates the instant of the anticipated AI-based prediction, while the gray label marks the actual moment of the spill. When a risk is detected, the AI system utilizes the 5G connection to activate a custom-built gripper to secure the liquid (represented in our tests by a water bottle), thereby ensuring operator safety and maintaining operational efficiency. In c), the left image shows the result when the PoC is inactive (i.e., the water bottle falls due to a sudden obstruction), while the right photo demonstrates how the proposed IIoT solution successfully prevents this outcome by accurately predicting the spill and activating the gripper in advance.
As an experimental demonstration of some of these research directions, Figure 2 shows the results of one of our PoC that leverages the interplay between AI and 5G NR for enhancing safety in a real-world UC belonging to the AD 4 family.
@@ IV-B Modeling Guidelines
In the following, we present modeling guidelines for industrial environments that adhere to the IIoT characterization given by this paper. This approach holds significant importance in establishing a cohesive framework for scientific endeavors aimed at addressing the research gaps previously highlighted. Indeed, the applicability of existing literature contributions to IIoT environments strongly depends on the level of realism of their underlying assumptions and characterizations. Moreover, the comparison between different solutions is possible only when they rely on common models.
@@
Industrial environments possess distinctive characteristics that render wireless propagation inherently unique and distinguishable from other radio channels. Typically, industrial premises exhibit larger dimensions and/or greater height compared to office or residential buildings. Furthermore, the size, density, and spatial distribution of industrial equipment can vary significantly from one case to another but, on average, constitute a highly dense scatterer environment. In light of these peculiarities, it becomes imperative to tailor channel models to suit industrial scenarios adequately. For instance, the 3GPP has proposed a path loss model specifically designed for indoor factories [^2] It is noteworthy that the term “factory” in 3GPP corresponds to the term “building” used in this article.[14]. However, more detailed wideband approaches can be obtained by leveraging (i) ray-tracing tools, as propagation in factories is likely to be more site-specific than propagation in usual residential or office environments, (ii) channel measurements, as a proper calibration method for ray-tracing tools, or (iii) the utilization of AI. Indeed, the latter is recently gaining attention to discern the intricate relationship between common propagation markers and the key properties of industrial environments [15].
We would also like to remark that experimental validation of models is inherently challenging in complex industrial systems due to limitations in current measurement techniques and data acquisition methods, which often struggle to capture the multidimensional interactions and dynamic behaviors necessary for comprehensive validation.
@@ V Conclusions
In this article, we presented a comprehensive characterization of the IIoT framework, a key enabler of the emerging fourth industrial revolution. Our focus included clarifying the diverse and sometimes conflicting contributions from entities like 3GPP and 5G-ACIA in terms of nomenclature, ADs, KPIs, and traffic types. After a careful literature analysis, we also outline the main IIoT requirements by harmonizing heterogeneous terminologies. Most importantly, we offered an overview of the main research topics and challenges within the IIoT paradigm according to the evolution of mobile radio networks, highlighting the potential for innovative industrial applications. Finally, we proposed modeling guidelines as a foundational framework for future research endeavors, addressing the identified research directions towards next-generation IIoT.